The CP of 20 eggs is equal to SP of 25 eggs. Find loss %

Grade:

2 Answers

Yash Baheti IIT Roorkee
9 years ago
Hi,

Let CP of each egg is 1Rs.
Then the selling price of 25 eggs is Rs. 20 (Cost price of 20 Eggs)
So SP of 1 egg is 0.8 Rs.

Loss is 20 % clearly

….….….….….….….….….….….….….
Debartha Paul
9 years ago
Let the CP of 1 egg be Rs x.
Therefore the CP of 20 eggs = Rs 20x
Therefore the CP of 25 eggs = Rs 25x
But CP of 20 eggs = SP of 25 eggs.
Therefore,
SP of 25 eggs= Rs 20x
Therefore net loss= Rs(25x-20x)= Rs 5x
Therefore net loss percentage= (5x/25x*100)% = 20%